Solve the equation in exact differentials: (y2-1)dx+(2xy+3y)dy=0


1
Expert's answer
2021-11-26T14:56:49-0500

"(y^2-1)dx+(2xy+3y)dy=0"

"Mdx+Ndy=0"

M=y2-1

N=2xy+3

"\frac{\partial{M}}{\partial{y}}=2y"

"\frac{\partial{N}}{\partial{x}}=2y"

"\therefore \frac{\partial{M}}{\partial{y}}=\frac{\partial{N}}{\partial{x}}" equation is exact

"\int Mdx=\int(y^2-1)dx=y^2x-x,(y=constant)"

"\int Ndy=\int(2xy+3y)dy=xy^2+\frac{3}{2}y^2,(x=constant)"

Hence the required solution is;

"xy^2-x+\frac{3}{2}y^2=C"

(rejecting the repeated term)